OpenStudy (anonymous):

2g-3h=0 3g-2h=5 Linear Combinations to solve the system of linear equations.

hero (hero):

Is linear combinations substitution or elimination? I dis-remember.

OpenStudy (anonymous):

Eliminiation

hero (hero):

2g-3h=0 3g-2h=5 g = 3h/2 g = (2h+5)/3 g = g 3h/2 = (2h+5)/3 2(2h+5) = 3(3h) 4h + 10 = 9h 10 = 9h - 4h 10 = 5h 10/5 =h 2 =h

hero (hero):

Good luck figuring that out

OpenStudy (anonymous):

Thanks So much!

hero (hero):

By the way, g= 3

OpenStudy (anonymous):

thanks